A flagpole 24 feet high is 10 feet away from the base of its anchor cable. How long is the anchor cable?
Answer by stanbon(75887)

A flagpole 24 feet high is 10 feet away from the base of its anchor cable. How long is the anchor cable?
------------------------------
Draw the picture:
You have a right triangle with height = 24' and base = 10'.
Find the diagonal (cable).
----
Use Pythagoras:
d^2 = 24^2 + 10^2
---
d = sqrt(676)
d = 26' (length of the cable)
===============================
Cheers,
Stan H.
===============================